### v2

Abstract
Real-time video segmentation is a promising feature for AI-assisted surgery, providing intraoperative guidance by identifying surgical tools and anatomical structures. However, deploying state-of-the-art segmentation models, such as SAM2, in real-time settings is computationally demanding, which makes it essential to balance frame rate and segmentation performance. In this study, we investigate the impact of frame rate on zero-shot surgical video segmentation, evaluating SAM2’s effectiveness across multiple frame sampling rates for cholecystectomy procedures. Surprisingly, our findings indicate that in conventional evaluation settings, frame rates as low as a single frame per second can outperform 25 FPS, as fewer frames smooth out segmentation inconsistencies. However, when assessed in a real-time streaming scenario, higher frame rates yield superior temporal coherence and stability, particularly for dynamic objects such as surgical graspers. Finally, we investigate human perception of real-time surgical video segmentation among professionals who work closely with such data and find that respondents consistently prefer high FPS segmentation mask overlays, reinforcing the importance of real-time evaluation in AI-assisted surgery.
Keywords:
AI-assisted surgery Computational efficiency in medical AI Real-time surgical video segmentation.
1 Introduction
Throughout history, surgical practice has been shaped by a series of transformative milestones that have fundamentally improved patient outcomes and expanded treatment possibilities [5]. Major advances include general anesthesia [18], antiseptic techniques [11], and imaging modalities like X-ray, CT, and MRI, which revolutionized surgical planning [6, 10]. The advent of minimally invasive techniques in the 20th century significantly reduced trauma, pain, and hospital stays [15] and more recently, the rise of computer-assisted and robotic surgery has increased surgical precision, reduced invasiveness, and enabled more complex procedures to be performed with greater accuracy [2]. Now, the next major frontier in surgery appears to be the integration of artificial intelligence (AI) at every stage of the surgical process [4].
Intraoperatively, AI-assisted robotic systems have the potential to enhance dexterity, accuracy, and consistency, enabling surgeons to operate with unprecedented control [23]. In particular, recent advancements in computer vision hold immense potential for real-time surgical video analysis, offering intraoperative guidance by identifying optimal dissection planes, highlighting high-risk areas, and marking safe zones for dissection [14]. Among state-of-the-art methods, video object segmentation (VOS) has recently emerged as a critical technology for surgical tool and organ segmentation [19]. Unlike traditional segmentation models that process frames independently, VOS models maintain spatial memory and track objects across frames, ensuring better consistency in real-time applications [16]. The SAM2 model, the state-of-the-art zero-shot segmentation framework, is particularly well-suited for surgical video analysis, as it effectively tracks and segments surgical instruments and anatomical structures without requiring task-specific fine-tuning [9, 17].
Figure 1: Two example images from the CholecSeg8k dataset with annotations for the gallbladder (purple), liver (light blue), and grasper (green).
Despite its potential, deploying such advanced AI models in real-time surgical applications comes with significant computational challenges [12]. VOS models such as SAM2 require powerful GPUs with large memories to maintain object tracking across frames. This limitation is particularly critical for prolonged surgeries, where the memory footprint of the tracking model grows continuously as more frames are processed [13]. Given the cost of high-performance GPUS, and the difficulty of obtaining them due to stock issues as well as export limitations, access to the necessary hardware may become a bottleneck for real-time AI adoption in surgery, especially for hospitals in countries that have difficulties in acquiring state-of-the-art GPUs [7]. To mitigate this issue, optimizing the ways in which frames are processed, for example by reducing the frame rate, may provide a way to lessen computational demands without compromising segmentation accuracy [22].
In this work, we evaluate the segmentation performance of the SAM2 model for cholecystectomy on three key targets: the gallbladder, liver, and surgical graspers. We conduct experiments across five frame-per-second (FPS) settings – 1, 10, 15, 20, and 25 FPS – to analyze the impact of frame rate on segmentation accuracy. Our findings reveal an intriguing phenomenon: when evaluated outside of real-time streaming scenarios, higher FPS settings may create the illusion of reduced performance. In extreme cases, our results show that the 1 FPS setting can outperform 25 FPS when assessed under conventional evaluation methods. However, when considering real-time streaming scenarios where predictions must remain stable across continuously arriving frames, 25 FPS setting yields superior performance, ensuring more consistent and temporally coherent segmentation. To complement these findings, we conducted a survey among surgeons, nurses, and machine learning engineers, assessing their perception of segmentation at different frame rates. Our findings indicate a strong preference for higher FPS segmentation mask overlays, reinforcing the importance of real-time evaluation in AI-assisted surgery.
2 Methodology
2.1 Model
In this study, we utilize SAM2.1 Hiera Large [17] (henceforth referred to as SAM2), a state-of-the-art zero-shot segmentation model, for real-time surgical video segmentation. SAM2 employs a powerful transformer-based architecture that ensures spatial and temporal consistency, making it particularly well suited for dynamic and complex environments such as surgical video analysis [3, 8, 21].
Several studies have demonstrated that SAM2 outperforms both traditional image-based segmentation models and other VOS frameworks, achieving state-of-the-art results in surgical video segmentation [24]. A key strength of SAM2 is its built-in tracking mechanism, which enhances temporal consistency across frames, reducing segmentation drift and improving overall coherence. This feature is particularly valuable in the operating room, where maintaining accurate and stable segmentation over time is crucial for real-time decision-making.
Table 1: Total number of frames used in each video for segmentation analysis, categorized by target objects (gallbladder, liver, and surgical grasper).
Objects Video 1 Video 12 Video 17 Video 20 Video 24 Video 35
Any 1,280 640 320 160 960 240
Gallbladder 992 624 320 160 686 240
Liver 1,280 640 320 160 960 240
Grasper 849 577 240 160 400 240
2.2 Data
For our experiments, we use the CholecSeg8k dataset, a frame-by-frame annotated dataset derived from Cholec80, a widely used surgical video dataset [1, 20]. Cholec80 consists of 80 cholecystectomy procedure videos recorded at 25 frames-per-second at the University Hospital of Strasbourg (Strasbourg, France) [20]. The CholecSeg8k dataset provides high-quality segmentation annotations for a subset of those videos, making it well-suited for evaluating real-time surgical video segmentation models [1].
In this study, we focus on the segmentation of two key anatomical structures—the gallbladder and liver—as well as a surgical instrument, the grasper. These targets are crucial for surgical scene understanding, as they provide insight into organ visibility and instrument interaction [14]. An example set of images from the dataset is provided in Figure 1.
Since the employed model is a zero-shot segmentation model, we do not need to partition the dataset into training and validation splits. Instead, we evaluate performance using six videos, some of which are divided into multiple segments. In total, we assess SAM2’s performance across five FPS settings – 1, 10, 15, 20, and 25 FPS – for 15 video segments. Table 1 shows the number of frames for each video that contain objects of each class.
2.3 Notation and Metrics
We represent a video as a sequence of $n$ number of RGB frames, denoted as $\mathsf{X}=[\texttt{X}_{1},\texttt{X}_{2},\ldots,\texttt{X}_{n}]$, where each frame $\texttt{X}\in\mathbb{R}^{3\times 854\times 480}$ corresponds to an image with three color channels (RGB) and a resolution of $854\times 480$ pixels.
To utilize the SAM2 model for video segmentation, an initial target object must be specified in the first frame, which will then be tracked throughout the sequence. In our approach, we define this target using a binary segmentation mask, represented as $\texttt{S}\in\{0,1\}^{854\times 480}$. This mask indicates the presence ($1$) or absence (00) of the object in each pixel.
Given an initial frame $\texttt{X}_{1}$ and mask $\texttt{S}_{1}$, SAM2 propagates segmentation across subsequent frames, predicting the $i$th frame as $g(\texttt{X}_{i})=\texttt{S}_{i}$. The mask $\texttt{S}_{i}$ depends on $\texttt{S}_{1}$ and all preceding frames with their inferred segmentations, ensuring temporal consistency in object tracking.
Intersection over Union (IoU). In order to evaluate the performance of predictions, we employ the IoU metric, which measures the overlap between the predicted segmentation mask S and the ground truth mask Y. Formally, IoU is defined as $\text{IoU}(\texttt{S},\texttt{Y})=\frac{|\texttt{S}\cap\texttt{Y}|}{|\texttt{S}\cup\texttt{Y}|}$.
Table 2: (Evaluation: Sampled frames) Average IoU scores for gallbladder, liver, and surgical grasper segmentation using the SAM2 model. Results are reported for different frame rates (1, 10, 15, 20, and 25 FPS), where all available frames in each FPS setting are evaluated. The best performance for each FPS setting per row is highlighted in bold.
Video Segment Gallbladder Liver Grasper
1 10 15 20 25 1 10 15 20 25 1 10 15 20 25
V1 S1 96.0 95.7 95.9 95.9 95.9 90.9 90.3 90.4 90.2 90.3 87.8 86.3 86.4 86.5 86.3
V1 S2 96.4 96.1 96.3 96.2 96.1 97.2 96.6 97.0 97.1 96.9 91.2 86.8 86.7 87.0 86.9
V1 S3 95.4 94.9 94.9 94.9 94.9 96.3 96.0 96.1 96.2 96.2 83.9 82.1 82.2 82.4 82.1
V1 S4 86.5 85.5 85.1 84.9 85.2 96.0 95.6 95.6 95.6 95.6 82.3 80.3 80.2 79.9 80.2
V1 S5 87.8 86.7 87.5 86.7 86.8 93.5 92.3 92.2 92.2 92.2 78.2 80.4 80.5 80.6 80.7
V12 S1 75.4 82.4 81.7 81.7 81.6 92.6 91.7 91.1 91.2 91.1 78.9 86.5 86.4 86.5 86.5
V12 S2 71.5 67.0 67.4 68.7 67.7 94.0 91.9 91.1 91.4 91.4 54.2 54.4 53.7 53.9 53.9
V12 S3 81.2 78.2 78.5 77.6 77.9 84.6 88.5 88.1 88.3 88.0 55.2 55.4 55.3 55.2 55.0
V17 S1 95.9 95.5 94.8 94.9 95.0 95.0 93.8 92.6 93.1 93.0 74.5 69.8 69.7 69.3 69.4
V17 S2 96.3 95.1 95.2 95.1 95.2 94.1 92.7 93.4 93.1 93.2 88.5 86.9 86.7 86.1 86.1
V20 S1 93.5 93.1 93.3 93.2 93.0 96.7 96.4 96.4 96.4 96.4 95.1 94.6 94.6 94.5 94.5
V24 S4 90.2 81.4 80.6 78.4 78.3 94.1 95.4 95.4 95.4 95.3 88.0 87.1 87.2 86.9 86.9
V35 S1 95.1 93.7 93.8 93.8 93.7 98.0 97.3 97.5 97.6 97.6 96.2 94.9 94.1 94.1 94.3
V35 S2 95.9 95.6 94.9 94.9 95.0 97.4 97.8 97.8 97.8 97.8 95.1 93.1 93.3 93.2 93.3
V35 S3 93.9 92.1 92.0 92.0 92.0 98.3 98.2 98.2 98.2 98.2 97.0 96.3 96.2 96.2 96.2
2.4 Evaluation
In this work, we evaluate the performance of SAM2 on surgical videos sampled in three settings: sampled frames, anchor frames, and real-time streaming frames.
Sampled frames. In this setting, we evaluate the IoU over sampled frames within the selected FPS setting. Formally, given a frame sequence $\mathsf{X}$ captured at 25 FPS, the evaluation is conducted over the subset $\mathsf{X}^{(f)}=[\texttt{X}_{1},\texttt{X}_{1+\Delta f},\texttt{X}_{1+2\Delta f},\dots]$, where $\Delta f=25/f$ determines the interval between selected frames. This method ensures that higher FPS settings use more frames for evaluation, but it introduces potential bias since different FPS settings involve different numbers of frames. For example, in a 10-second video recorded at 25 FPS, there are $10\times 25=250$ total images, whereas at 1 FPS, there are only $10\times 1=10$ images. This means that higher FPS settings provide more temporal details, which can influence the evaluation results. To mitigate this bias, we introduce the anchor frames setting.
Anchor frames. In this setting, we standardize the number of frames used for evaluation across different FPS settings. Specifically, we define the anchor frame set as $\mathsf{X}_{\text{Anchor}}=[\texttt{X}_{1},\texttt{X}_{26},\texttt{X}_{51},\dots]$, where only the first frame of each second is used, ensuring that all FPS settings are evaluated on the same number of frames. This approach eliminates potential biases introduced by varying frame counts and provides a fairer comparison across different FPS configurations.
Real-time streaming frames. This setting simulates a real-time streaming scenario where predictions from lower FPS settings must persist across intermediate frames. Regardless of the selected FPS setting, evaluation is conducted on all 25 frames per second. Formally, for an FPS $f$, predictions $\texttt{S}_{t}$ at sampled frames are propagated forward across intermediate frames until the next sampled frame appears. That is, for an FPS setting of $f$, a prediction made at $\texttt{X}_{t}$ remains unchanged for frames $\texttt{X}_{t+1},\dots,\texttt{X}_{t+\Delta f-1}$ until the next prediction update occurs at $\texttt{X}_{t+\Delta f}$. This setup assesses the practical impact of frame rate on segmentation performance in real-time applications, where predictions must remain valid between updates to prevent flickering.
Table 3: (Evaluation: Anchor frames) Average IoU scores for gallbladder, liver, and surgical grasper segmentation using the SAM2 model when evaluated only on anchor frames (the first frame of each second). The best performance for each FPS setting per row is highlighted in bold.
Video Segment Gallbladder Liver Grasper
1 10 15 20 25 1 10 15 20 25 1 10 15 20 25
V1 S1 96.3 96.4 96.4 96.5 96.5 90.9 91.1 91.0 90.9 90.6 87.8 87.9 88.1 88.0 87.9
V1 S2 97.2 97.2 97.2 97.2 97.2 97.2 97.2 97.2 97.2 97.2 91.2 90.5 90.2 90.5 90.6
V1 S3 95.4 95.6 95.6 95.5 95.4 96.3 96.4 96.4 96.5 96.5 96.2 96.2 96.2 96.2 96.2
V1 S4 94.1 95.7 95.8 95.8 95.7 96.0 95.9 96.0 95.9 95.9 75.4 77.8 77.5 77.7 78.1
V1 S5 93.5 93.3 93.4 93.4 93.4 93.5 93.3 93.4 93.4 93.4 82.3 82.4 82.3 82.1 82.1
V12 S1 95.0 95.0 94.9 94.9 95.0 92.6 92.8 92.2 92.2 92.2 83.5 83.4 83.5 83.4 83.2
V12 S2 96.7 96.7 96.6 96.6 96.7 94.0 94.1 94.0 94.0 93.9 95.0 95.0 94.9 94.9 95.0
V12 S3 94.1 94.0 94.0 93.9 94.1 84.6 88.8 87.0 88.9 86.9 89.4 89.3 89.3 89.4 89.1
V17 S1 95.9 95.8 95.8 95.8 95.9 95.0 95.1 95.1 95.1 93.8 95.2 95.3 95.3 95.2 95.2
V17 S2 96.0 96.0 96.1 96.1 96.0 94.1 94.1 94.1 94.1 94.1 95.3 95.3 95.3 95.2 95.1
V20 S1 96.8 96.8 96.8 96.8 96.8 96.7 96.8 96.8 96.8 96.8 95.3 95.3 95.3 95.2 95.1
V24 S4 90.2 80.2 79.9 76.1 76.7 94.1 95.7 95.8 95.8 95.7 95.7 95.8 95.8 95.7 95.7
V35 S1 98.3 98.5 98.5 98.5 98.5 98.0 98.2 98.2 98.2 98.2 96.8 96.8 96.8 96.8 97.0
V35 S2 98.0 98.0 98.0 98.0 98.0 97.4 98.0 98.0 98.0 98.0 95.2 95.3 95.3 95.2 95.1
V35 S3 97.4 98.0 98.0 98.0 98.0 98.3 98.5 98.5 98.5 98.5 96.8 96.8 96.8 96.8 97.0
3 Experimental Results
In this section, we present the evaluation results of the SAM2 model across the three evaluation strategies highlighted in Section 2.4: (1) sampled frames, (2) anchor frames, and (3) real-time streaming frames. Results for these settings can be found in Table 2, Table 3, and Table 4, respectively.
Sampled frames. This setting assesses segmentation accuracy by computing the IoU for the sampled frames available in a given FPS setting. Surprisingly, in most cases, the 1 FPS setting outperforms the 25 FPS setting (see Table 2). This counterintuitive result arises because lower FPS evaluates performance on significantly fewer frames. By selecting only one frame per second, this setting inherently smooths out segmentation inconsistencies, making the results appear more stable and artificially inflating IoU scores. Higher FPS settings, on the other hand, introduce more frequent updates, which expose even minor segmentation variations. These variations, while not necessarily indicative of poor performance, result in a slight reduction in IoU scores. Our observations suggest that this evaluation setting may not be the most reliable method for assessing real-world segmentation stability, as it can mask critical segmentation inconsistencies during operations that become apparent at higher frame rates.
Table 4: (Evaluation: Real-time streaming) Average IoU scores for gallbladder, liver, and surgical grasper segmentation using the SAM2 model in a real-time streaming scenario, where lower FPS predictions persist across intermediate frames. The best performance for each FPS setting per row is highlighted in bold.
Video Segment Gallbladder Liver Grasper
1 10 15 20 25 1 10 15 20 25 1 10 15 20 25
V1 S1 79.8 94.1 95.1 95.5 95.9 81.8 89.4 89.9 89.9 90.3 30.8 35.9 39.7 42.2 74.4
V1 S2 88.5 95.5 95.8 96.0 96.1 90.9 95.9 96.4 96.6 96.9 57.5 83.3 84.6 85.9 86.9
V1 S3 84.2 94.0 94.5 94.7 94.9 83.3 94.8 95.5 95.9 96.2 63.5 84.4 85.5 86.0 86.3
V1 S4 67.7 82.7 83.9 84.2 85.2 79.9 94.0 94.8 95.2 95.6 37.7 73.5 77.0 79.4 82.1
V1 S5 61.8 84.1 86.3 86.2 86.8 81.3 91.5 91.9 92.0 92.2 37.4 75.4 77.8 79.2 80.2
V12 S1 62.1 80.0 81.6 81.5 81.6 79.1 90.5 90.8 91.0 91.1 34.8 73.9 77.8 79.2 80.7
V12 S2 57.8 67.0 67.0 68.8 67.7 83.7 90.8 91.2 91.4 91.4 41.5 52.6 53.1 53.8 53.9
V12 S3 59.5 76.9 77.4 77.8 77.9 70.4 86.8 87.5 87.8 88.0 38.7 80.5 83.5 84.8 86.5
V17 S1 66.2 92.0 93.5 94.4 95.0 69.3 90.4 91.6 92.4 93.0 30.9 64.0 66.4 67.7 69.4
V17 S2 78.6 93.0 94.1 94.6 95.2 74.8 91.4 92.3 92.8 93.2 34.7 75.9 80.1 83.3 86.1
V20 S1 83.5 92.1 92.7 93.0 93.0 89.4 95.9 96.2 96.3 96.4 79.2 92.7 93.7 94.2 94.5
V24 S4 45.1 81.7 83.6 85.0 86.3 74.2 90.2 91.3 91.9 92.4 86.9 83.2 80.6 75.9 31.2
V35 S1 56.3 85.6 89.9 91.8 93.7 83.6 95.7 96.7 97.1 97.6 29.8 77.3 85.6 90.4 94.3
V35 S2 56.2 90.3 92.7 93.9 95.0 73.6 95.5 96.9 97.5 97.8 25.2 78.8 86.5 89.9 93.3
V35 S3 69.0 89.4 90.7 91.3 92.0 86.6 97.2 97.7 97.9 98.2 45.8 87.8 91.7 94.1 96.2
Anchor frames. To address potential biases introduced by varying frame counts across FPS settings, the anchor frames setting evaluates segmentation only on the first frame of each second. As shown in Table 3, this setting reveals no apparent superior FPS choice, with segmentation performance remaining relatively similar across all frame rates.
Real-time streaming frames. This setting simulates real-world deployment by requiring lower FPS predictions to persist across intermediate frames to prevent flickering. This evaluation reflects practical constraints in real-time applications, where model outputs must remain stable between updates. As illustrated in Table 4, segmentation accuracy consistently favors 25 FPS, particularly for the surgical grasper, which is the object that moves by far the most in surgical operations. Unlike traditional evaluation settings, real-time evaluation clearly demonstrates that higher FPS settings are crucial for maintaining segmentation consistency and preventing temporal drift.
3.1 Human Perception of Annotation FPS in Surgical Videos
Following our experiments, we conducted a survey to evaluate the human perception of segmentation frame rates in surgical videos. We asked participants, including surgeons, nurses, and machine learning engineers, to compare videos with segmentation maps generated at different frame rates (e.g., 25v22, 25v20, 25v15) and indicate whether they preferred the higher frame rate, the lower frame rate, or had no preference. Participants were only provided with the segmented video, and did not know the underlying segmentation frame rate. Our goal in this survey is to assess how variations in segmentation FPS impact perceived quality and whether lower FPS affects visual perception.
[…]
Figure 2: Illustration of how surgeons, nurses, and engineers perceive differing segmentation frame rates in 25 FPS surgical videos. Each grouped set of bars represents a frame rate comparison between two videos, with responses categorized into three choices: preferring the first video/FPS (left bar), preferring the second video/FPS (middle bar), or no preference/either (right bar).
4 Conclusion and Future Perspectives
In this work, we revisited the impact of frame rate sampling on real-time zero-shot surgical video segmentation by evaluating the performance of the SAM2 model across multiple frame sampling rates. We discovered that, under conventional evaluation settings, lower frame rates tend to yield higher segmentation scores than highper FPS settings due to the smoothing effect of fewer segmentation updates. However, experiments conducted under real-time streaming conditions demonstrated that higher frame rates nevertheless achieve the highest segmentation scores, ensuring better temporal coherence and stability.
To further investigate the perceptual impact of segmentation frame rates, we conducted a survey involving surgeons, nurses, and machine learning engineers. The results of this survey revealed that as the gap between segmentation mask overlays increased, participants showed a stronger preference for higher FPS settings, reinforcing the necessity of real-time evaluation in AI-assisted surgery.
With modern laparoscopic systems supporting frame rates up to 120 FPS, the impact of frame rate on AI-assisted segmentation remains an open question. At lower frame rates such as 25, humans can easily identify flickering and segmentation lags, since human visual perception is generally sensitive to motion inconsistencies and operates optimally between 30 to 60 FPS. However, at higher frame rates, segmenting every frame in real time may not be necessary, as differences between consecutive frames become less perceptible. Future work should explore adaptive frame rate processing for high FPS surgical videos, selectively applying segmentation based on scene dynamics to optimize computational efficiency without compromising quality.